OpenStudy (anonymous):

could someone help me and explain me how to solve this: t+1/2t-1 = 5/7 thanks!

OpenStudy (anonymous):

If you have the equation that T + 1/ 2t -1 = 5/7, you could cross multiply to get 5(2t-1) = 7(T +1) 10t - 5 = 7T + 7 3T = 12 t = 4